OpenStudy (anonymous):

Bernie counted the cars and motorcycles that passed his house in 45 minutes. He counted 20 vehicles and 72 tires. How many cars passed his house? 8 12 16 20

OpenStudy (campbell_st):

let x = cars and y = motorcycles.. now the equation for the number of cars is x + y = 20 the number of tires is 4x + 2y = 72 ( a car has 4 tires, a motorcycle has 2 tires) so you have to solve the equations simulataneously. the easiest way is to use substitution and look to eliminate y so from the 1st equation x + y= 20 rewrite it with y as the subject y = 20 - x now substitute it into the 2nd equation 4x + 2y = 72 or 4x +2(20 - x) = 72 now you can sole for x, the number of cars

OpenStudy (anonymous):

so then it would be 4 motorcycles and 16 cars?

OpenStudy (campbell_st):

correct